If,we have the following function F(x) =(2x-2-x) /(2x+2-x)
Find f-1(x). 
 
           

Grade:12

1 Answers

Susmita
5 years ago
Write f(x)=y.
So 
y=
f(x)=\frac{2^x -\frac{1}{2^x}}{2^x+\frac{1}{2^x}}
=\frac{4^x -1}{4^x+1}
Swap x and y.
Or,x=\frac{4^y -1}{4^y +1}
Or,\frac{x+1}{x-1} =\frac{(4^y -1)+(4^y+1)}{(4^y-1)-(4^y +1)}
Or,\frac{x+1}{x-1} =\frac{(2*4^y)}{(-2)}
Or,\frac{x+1}{x-1} =-4^y
Take logarith on both side.
Or,\log \frac{x+1}{x-1} =-y\log 4
Find y and write y=f-1(x).
Or,y=\log ( \frac{x-1}{x+1} -4)=f^-1(x)
Hope this helps.